A machine manufactures spare parts at the rate of 20, 000 per month. A second machine uses those spare parts at the rate 50 00 per month and remainder put into stock. It cost ₦100, 000 to set up the machine the company establish their stock hold costs 20% per annum of the average stock value each parts costs ₦250 to make. Required What batch size should be produced on the first machine and what frequency.
1
Expert's answer
2020-02-12T07:33:58-0500

To comply with all conditions the first machine should produce 20000 units. But we can't produce this amount each month as there is restriction on stock hold costs, so the first machine should produce 20000units each 4 months. The point is that we need to have stock not less than 20% for the second machine. From the first one we get 250*20000 = 5 000 000. But for the set up 5000 000 - 100 000 = 4 900 000. At the beginning it was possible to produce 20000 each 5 month as 5000 000 / 4 = 1250 000 (5000*250). But after subtracting the set up value it became less, so we need to get new parts each 4 months.